OpenStudy (anonymous):

Please help me, will medal & fan! What are the intercepts of 3x − 9y = -27? (-3, 0) and (0, 9) (3, 0) and (0, 9) (3, 0) and (0, 6) (-9, 0) and (0, 3)

OpenStudy (anonymous):

@Conqueror

OpenStudy (igreen):

Just plug in '0' for 'x' and 'y'. 3x - 9y = -27 Plug in 0 for 'x': 3(0) - 9y = -27 -9y = -27 Divide -9 to both sides, what's -27 / -9?

OpenStudy (anonymous):

Multiply x values times 3 and the y values by 9 Then subtract y from x and see if they equal -27...

OpenStudy (igreen):

Can you divide -27 / -9? @itsdaniela

OpenStudy (anonymous):

Its 3 @Conqueror @Heypips @iGreen

OpenStudy (igreen):

Yes, so the y-intercept is (0, 3)

OpenStudy (anonymous):

Yes

OpenStudy (igreen):

Now we can plug in '0' for 'y' to find the x-intercept. 3x - 9y = -27 3x - 9(0) = -27 3x = -27 Divide 3 to both sides, what's -27 / 3?

OpenStudy (anonymous):

-9 @iGreen

OpenStudy (igreen):

Yes, so the x-intercept is (-9, 0).

OpenStudy (igreen):

So we have: (-9, 0) and (0, 3)

OpenStudy (anonymous):

Thank you so much! :) @iGreen

OpenStudy (igreen):

No problem.

OpenStudy (anonymous):

Can you help me with this problem? What are the intercepts of y = -0.75x + 3.0?

OpenStudy (anonymous):

please? @iGreen

OpenStudy (igreen):

Sure, but ask it in a new question.

OpenStudy (anonymous):

okay.